OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (anonymous):

im not sure what i am doing.

OpenStudy (anonymous):

where is f(x) which you created in problem number 2 ? and i highly doubt that its question from test?

OpenStudy (anonymous):

i used F(x)=5x+2 but im not even sure if itll work

OpenStudy (anonymous):

I Seriously need a grade for my flvs online i just dont understand this .

OpenStudy (anonymous):

lets say we assinged number n to x so that h(x) would become n = > h(x) = n then f(h(x)) = f(n) = 5n + 2 h(f(x)) = h(5x+2) = h(5x)+2 = 5h(x)+2 = 5n +2 so they both are equal

OpenStudy (anonymous):

did i make any sense?

OpenStudy (anonymous):

somewhat is that how i will explain it in my response?

OpenStudy (anonymous):

does that seem to be convincing for you?, if you understood what i said, you will be able to write it in your way, nonetheless tests test to which point you can elaborate the question in hand

OpenStudy (anonymous):

thank you i understand !!